# Load Forecasting for Households and Energy Communities: Are Deep Learning Models Worth the Effort?

### v3

@@ Abstract
keywords:
Load Forecasting , Energy Communities , Deep Learning , Transformer , xLSTM , Transfer Learning
journal: Applied Energy
[…]
@@ 1 Introduction
In practice, MPC requires accurate forecasts of exogenous variables such as electrical power demand, heat load, or photovoltaic (PV) production. Many recent studies have been affiliated with heat load forecasting using data-driven or physics-based methods [5, 6, 7]. However, the highest relevance for ECs lies in an accurate prediction of the electric power. A meaningful distinction can be drawn between long-term and short-term power forecasting: the former is crucial for strategic planning and infrastructure development [8], whereas the latter is essential for real-time optimization of flexible energy assets [9, 10].
@@
Several recent studies highlight the superior performance of Transformer-based models over LSTMs in the context of short-term load forecasting. For example, Semmelmann et al. [13] compared the forecasting errors of LSTMs and Transformers for an EC consisting of 21 households and concluded that Transformer models outperform LSTMs. Other independently conducted studies [14, 15, 16] all came to similar conclusions when evaluating single highly aggregated load profiles representing entire cities and regions. L’Heureux et al. [17] additionally showed that a Transformer-based model outperforms a state-of-the-art recurrent model when tested on 20 highly aggregated load profiles from an energy utility company.
@@ 2 Methods
In accordance with Fig. 1, different types of data were used and curated to evaluate various models regarding their load forecasting capabilities. Typical input features for short-term load forecasting were derived, including datetime information, weather conditions, and historical load data, which corresponds to other load forecasting studies [23, 8, 13, 6].
@@ 4 Conclusions
Lukas Moosbrugger: Writing – Original Draft, Visualization, Software, Validation, Methodology, Data curation, Formal analysis, Conceptualization. Valentin Seiler: Writing – Review & Editing, Visualization, Validation, Methodology, Conceptualization. Philipp Wohlgenannt: Writing – Review & Editing, Software, Validation, Methodology, Conceptualization. Sebastian Hegenbart: Writing - Review & Editing, Methodology, Conceptualization. Sashko Ristov: Writing - Review & Editing, Methodology, Conceptualization. Elias Eder: Writing - Review & Editing, Visualization, Validation, Methodology, Conceptualization. Peter Kepplinger: Writing – Review & Editing, Conceptualization, Supervision, Resources, Project administration, Funding acquisition.
@@ Appendix C Model Configuration
0.2k
Blocks = 1
Heads = 1
Features = 1
sLSTM at 0
@@
80k Blocks = 4 Heads = 8 Features = 40 sLSTM at 1

### v4

@@ 1 Introduction
In practice, MPC requires accurate forecasts of exogenous variables such as electrical power demand, heat load, or photovoltaic (PV) production. Many recent studies have been affiliated with heat load forecasting using data-driven or physics-based methods [5, 6, 7]. However, the highest relevance for ECs lies in an accurate prediction of the electric power. A meaningful distinction can be drawn between long-term and short-term power forecasting: the former is crucial for strategic planning and infrastructure development [8, 9], whereas the latter is essential for real-time optimization of flexible energy assets [10, 11].
@@
Several recent studies highlight the superior performance of Transformer-based models over LSTMs in the context of short-term load forecasting. For example, Semmelmann et al. [14] compared the forecasting errors of LSTMs and Transformers for an EC consisting of 21 households and concluded that Transformer models outperform LSTMs. Other independently conducted studies [15, 16, 17, 18] all came to similar conclusions when evaluating single highly aggregated load profiles representing entire cities and regions. L’Heureux et al. [19] additionally showed that a Transformer-based model outperforms a state-of-the-art recurrent model when tested on 20 highly aggregated load profiles from an energy utility company.
@@ 2 Methods
In accordance with Fig. 1, different types of data were used and curated to evaluate various models regarding their load forecasting capabilities. Typical input features for short-term load forecasting were derived, including datetime information, weather conditions, and historical load data, which corresponds to other load forecasting studies [25, 8, 14, 6, 18, 27].
@@ 4 Conclusions
Lukas Moosbrugger: Writing – Original Draft, Visualization, Software, Validation, Methodology, Data curation, Formal analysis, Conceptualization. Valentin Seiler: Writing – Review & Editing, Visualization, Validation, Methodology, Conceptualization. Philipp Wohlgenannt: Writing – Review & Editing, Software, Validation, Methodology, Conceptualization. Sebastian Hegenbart: Writing - Review & Editing, Validation, Methodology, Conceptualization. Sashko Ristov: Writing - Review & Editing, Validation, Methodology, Conceptualization. Elias Eder: Writing - Review & Editing, Visualization, Validation, Methodology, Conceptualization. Peter Kepplinger: Writing – Review & Editing, Validation, Conceptualization, Supervision, Resources, Project administration, Funding acquisition.
@@ Appendix C Model Configuration
0.2k Blocks = 1 Heads = 1 Features = 1 sLSTM at 0
@@
80k
Blocks = 4
Heads = 8
Features = 40
sLSTM at 1